# Differential Gated Self-Attention

### v1

@@ 5.2 Language Classification
In our language classification experiments, we progressively evaluate our model on tasks of increasing difficulty by the number of target classes. We begin with binary classification involving 2 classes, advance to a 4-class setting, and finally test the model’s capacity on a more complex 20-class classification task. This gradual increase in the number of classes allows us to systematically assess the model’s performance and robustness across varying levels of classification complexity.
Datasets.
Each dataset comes with pre-split training and test sets and we derive a validation split by dividing the original training set into 80% for training and 20% for validation. The Rotten Tomatoes  (Pang and Lee, 2005) and the IMDB  (Maas et al., 2011) movie reviews datasets are for binary sentiment classification. The AG News dataset (Del Corso et al., 2005), derives from AG’s corpus of news articles, focused on four major news categories. The 20 Newsgroups dataset (Mitchell, 1997) contains newsgroup messages organized into 20 distinct categories.
Table 2: Classification accuracy (mean ± std) on language benchmarks. Averaged over 5 seeds.
Rotten Tomatoes (%) IMDB (%) AGNews (%) 20 Newsgroups (%)
#Classes 2 2 4 20
#Size 10,000 50,000 127,600 18,000
Transformer $72.18\pm 0.51$ $85.34\pm 0.22$ $91.61\pm 0.08$ $51.54\pm 0.49$
DIFF Transformer
@@
   ×$\times 2$ $73.99\pm 0.91$ $86.41\pm 0.14$ $91.85\pm 0.23$ $46.38\pm 0.96$
   ×$\times 4$ $73.06\pm 0.7$ $86.49\pm 0.29$ $92.03\pm 0.2$ $47.38\pm 2.08$
   ×$\times 16/3$ $73.19\pm 0.85$ $86.43\pm 0.16$ $92.00\pm 0.19$ $49.99\pm 2.80$
@@
   ×$\times 2$ $74.26\pm 0.97$ $86.47\pm 0.17$ $\mathbf{92.25\pm 0.14}$ $60.32\pm 0.49$
   ×$\times 4$ $73.83\pm 1.15$ $\mathbf{86.97\pm 0.43}$ $91.89\pm 0.14$ $59.07\pm 0.48$
   ×$\times 16/3$ $\mathbf{74.40\pm 0.42}$ $86.53\pm 0.33$ $91.98\pm 0.27$ $\mathbf{63.53\pm 0.87}$
@@
We compare our DGT model’s performance to both the vanilla Transformer and the Differential Transformer under default configurations. For a fair comparison, we evaluate Differential Transformer variants with feed-forward modules expanded by factors of 2, 4, and 16/3 relative to the output dimension $Y$, applying SwiGLU before projecting back to the base dimension, where the 16/3 expansion matches the original Differential Transformer setup. We then apply these identical expansion factors and activation sequence to our DGT model, aligning experimental conditions across all architectures to isolate the impact of our approach across multiple benchmark tasks and comprehensive evaluation protocols.
@@ Results.
Our results in Table 2 indicate that incorporating Differential Gated Attention improves prediction accuracy in language classification tasks. Specifically, we observe that the DGT model’s performance increases as the complexity of the classification task grows, particularly with a higher number of target classes. While the model achieves a modest improvement of 1-2% on the Rotten Tomatoes, IMDB, and AG News datasets, it shows a substantial gain of 12-14% on the more challenging 20 Newsgroups task. This trend suggests that the model not only scales well with increased task difficulty but also benefits from the richer structure and diversity present in fine-grained classification settings.
@@
Figure 2: Attention-rollout heatmaps for DGViT vs ViT on CIFAR-100 (top two rows) and SVHN (bottom two rows). Bright yellow indicates high attention weights, whereas deep blue denotes low or suppressed attention. While ViT’s attention is broadly distributed, assigning high weights to both the target and surrounding background pixels, DGViT exhibits greater discriminative focus.
Vision (DGViT vs ViT).
Fig. 2 shows attention-rollout heatmaps for DGViT versus the standard ViT on CIFAR-100 and SVHN. In each row, the left panel is the input image, the middle, DGViT’s fused attention, and the right, ViT’s attention. DGViT concentrates excitatory weights (bright yellow) on semantically salient structures - e.g., mushroom contours and digit strokes - while assigning inhibitory weights (deep blue) to background clutter, highlighting its fine-grained boundary delineation and illustrating selective focus on semantically salient regions.
@@
(b) DIFF-Transformer model attention
@@
Figure 3: Attention-rollout for a 20 Newsgroups sample. Bright yellow indicates strong positive attention, and deep blue denotes inhibitory weights. (a) The standard Transformer baseline spreads its attention diffusely across both function words and content tokens, failing to isolate the class-defining terms. (b) The Differential Transformer partially suppresses common-mode noise, but still assigns non-negligible weight to other classes too. (c) Our DGT model sharply concentrates on the key parts of the input while attenuating filler words. The focused, input-dependent gating aligns attention with the correct comp.windows.x label and filters out background noise.
@@
Fig. 3 presents attention-rollout for (a) Transformer, (b) Differential Transformer, and (c) our DGT on a 20 Newsgroups sample, input text: "Most graphics systems I have seen have drawing routines that also specify a color for drawing, like Drawpoint(x,y,color) or Drawline(x1,y1,x2,y2,color) or Fillrectangle(x1,y1,x2,y2,color)…With X, I…", label: comp.windows.x. The vanilla Transformer disperses its weights broadly over both common function words and various candidate labels, failing to single out the true class while the Differential Transformer partly suppresses common-mode noise yet still spreads attention to competing classes, diluting its focus on the correct one. DGT’s input-dependent gating, in contrast, concentrates focus on key class-indicative words, filtering out irrelevant text and yielding a more interpretable, discriminative attention pattern, increasing attention mass on ground-truth and aligning sharply with comp.windows.x.
@@ 6 Discussion
Our results demonstrate that embedding an input-conditioned lateral inhibition mechanism into self-attention yields substantial gains in noise resilience and generalization across both vision and language tasks. M-DGSA consistently sharpens attention maps, assigning strong excitatory weights to semantically relevant regions while suppressing background noise, without requiring any task-specific tuning. Its capacity to delineate fine boundaries suggests applications in safety-critical domains - such as lane-marking detection in autonomous driving or edge delineation in medical imaging—where robust, fine-grained attention is paramount.
Despite these benefits, several limitations invite further investigation. First, we have evaluated on mid-scale benchmarks with synthetic noise, extending to real-world distortions (e.g., motion blur, occlusions) and large-scale tasks (e.g., ImageNet classification) may reveal additional strengths or failure modes. Second, we confined M-DGSA to self-attention, integrating lateral inhibition into cross-attention, such as in vision-language models for VQA or image captioning, could enable dynamic, question-conditioned suppression of irrelevant image regions, further enhancing multimodal alignment.
For future work, we see several promising directions: extending lateral-inhibition gating to cross-attention and multimodal models, scaling to large-scale benchmarks and real-world corruption regimes, and developing hardware-efficient implementations that exploit the induced attention sparsity. We also anticipate that a deeper theoretical understanding of learned inhibition patterns will inform principled design and initialization of future attention mechanisms.
@@ 7 Conclusion
We have introduced Multihead Differential Gated Self-Attention (M-DGSA), a lightweight yet powerful extension of multihead self-attention that embeds per-head input-dependent lateral inhibition into the dual-softmax framework. Leveraging dynamic, context-sensitive gating, M-DGSA achieves sharper, more selective attention maps, delivering enhanced noise resilience and improved cross-domain performance on both vision and language benchmarks. Importantly, it integrates seamlessly into existing Transformer and ViT architectures with negligible overhead, making it immediately applicable to a wide range of tasks.
Beyond empirical improvements, M-DGSA advances the interpretability of attention mechanisms: its gating patterns mirror biological contrast-enhancement, providing clear, roll-out visualizations of how models amplify salient features and quench noise. We believe this biologically inspired paradigm marks a significant step toward more robust, trustworthy attention models for next-generation deep learning.
@@ A.1 Detailed description of the used datasets
AG News. (Del Corso et al., 2005) Derived from AG’s corpus of news articles, the AG News dataset comprises four categories—World, Sports, Business, and Science/Technology—by combining each article’s title and description and selecting the four largest classes. It contains $120,000$ training samples ($30,000$ per category) and $7,600$ test samples ($1,900$ per category).
20 Newsgroups. (Mitchell, 1997) $\sim 18,000$ posts organized into 20 distinct newsgroups, each corresponding to a specific topic such as computer hardware, recreational activities, politics, science, or religion. We adopt the canonical split of $11,314$ training and $7,532$ test posts, yielding on average about $566$ training and $377$ test samples per class.
@@
Street View House Numbers (SVHN). (Netzer et al., 2011) A large-scale, real-world image dataset comprising over 600,000 32×32-pixel digit crops extracted from Google Street View house numbers. It includes 10 classes (digits 0–9): 73,257 training images, 26,032 test images, and 531,131 additional “extra” samples of comparatively easier digits. By capturing digits in varied natural scenes—with diverse backgrounds, lighting conditions, and occlusions—SVHN poses a substantially more challenging recognition task than MNIST.
@@ A.2 Dataset Splits
For each language dataset, AG News, IMDB, 20 Newsgroups, and Rotten Tomatoes, we further split the training fold into $80\%$ train and $20\%$ validation (stratified by class). All final results are reported on the official test set. For all vision benchmarks, CIFAR-10, CIFAR-100, MNIST, Fashion-MNIST, and SVHN, we use the standard train/test partitions provided by each dataset and no further hold-out split was applied.
@@ A.3.2 Language Benchmarks
All text-classification experiments use the hyperparameters listed in Table 4. For the Rotten Tomatoes dataset, we instead employ 4 layers, a learning rate of $45e-$ 4, and 500 warmup steps.
Table 4: Hyperparameters for text datasets.
@@ A.4 Compute & Runtime
All experiments ran in PyTorch on a single NVIDIA A10G (24 GB VRAM).
  • Vision: $\sim 0.5$min/epoch on CIFAR-10/100, $\sim 0.6$ min on FashionMNIST, $\sim 0.76$ min on SVHN.
  • Language: $\sim 0.3$min/epoch on Rotten Tomatoes, $\sim 0.39$ min on IMDB, $\sim 0.57$ min on AG News, $\sim 0.5$ min on 20NewsGroups
Table 5: Number of learnable parameters per epoch for vision benchmarks.
CIFAR-10 CIFAR-100 FashionMNIST SVHN
   DGViT
   proj. ×$\times 2$ $3.77$M $3.8$M $3.77$M $3.77$M
   proj. ×$\times 4$ $5.35$M $5.36$M $5.35$M $5.35$M
Table 6: Number of learnable parameters per epoch for language benchmarks.
@@
   DGT
   proj. ×$\times 2$ $3.8$M $25.7$M $13.2$M $12$M
   proj. ×$\times 4$ $4.6$M $26.9$M $14.4$M $12.8$M
   proj. ×$\times 16/3$ $5.1$M $27.7$M $15.2$M $13.4$M

### v2

@@ 5.2 Language Classification
For language classification, we design a progression of tasks that increase in difficulty with respect to the number of target classes and the diversity of linguistic structure. The evaluation begins with binary sentiment classification (Rotten Tomatoes (Pang and Lee, 2005), IMDB (Maas et al., 2011)), progresses to four-way news categorization task (AG News (Del Corso et al., 2005)), and extends to the 20-class 20 Newsgroups benchmark  (Mitchell, 1997), which demands fine-grained discrimination across diverse categories. To complement this controlled scaling, we additionally evaluate on MNLI (Williams et al., 2018), a large-scale natural language inference benchmark spanning multiple genres, which serves as a rigorous test of generalization under heterogeneous conditions. Together, this suite of datasets provides a principled framework for assessing both the performance and robustness of our approach across varying levels of classification complexity.
Table 3: Classification accuracy (%) (mean ± std) on language benchmarks. Averaged over 5 seeds.
Rotten Tom. IMDB MNLI AGNews 20 Newsg.
#Classes 2 2 3 4 20
#Size 10,000 50,000 433,000 127,600 18,000
Transformer $72.18\pm 0.51$ $85.34\pm 0.22$ $55.03\pm 0.01$ $91.61\pm 0.08$ $51.54\pm 0.49$
DT
@@
   $\times 2$ $73.99\pm 0.91$ $86.41\pm 0.14$ $57.00\pm 0.004$ $91.85\pm 0.23$ $46.38\pm 0.96$
   $\times 4$ $73.06\pm 0.70$ $86.49\pm 0.29$ $56.11\pm 0.01$ $92.03\pm 0.20$ $47.38\pm 2.08$
   $\times 16/3$ $73.19\pm 0.85$ $86.43\pm 0.16$ $56.90\pm 0.003$ $92.00\pm 0.19$ $49.99\pm 2.80$
@@
   $\times 2$ $74.26\pm 0.97$ $86.47\pm 0.17$ $57.60\pm 0.004$ $\mathbf{92.25\pm 0.14}$ $60.32\pm 0.49$
   $\times 4$ $73.83\pm 1.15$ $\mathbf{86.97\pm 0.04}$ $\mathbf{58.07\pm 0.001}$ $91.89\pm 0.14$ $59.07\pm 0.48$
   $\times 16/3$ $\mathbf{74.40\pm 0.42}$ $86.53\pm 0.33$ $57.26\pm 0.004$ $91.98\pm 0.27$ $\mathbf{63.53\pm 0.87}$
@@
We compare our DGT model’s performance to both the vanilla Transformer and the Differential Transformer (DT) under default configurations. To ensure fairness and to isolate the contribution of the attention mechanism, we adopt identical experimental conditions for DT and DGT: the feed-forward module expands the hidden dimension by factors of 2, 4, and 16/3 relative to the output dimension $Y$, applying SwiGLU before projecting back to the base dimension. This alignment guarantees that performance differences can be attributed to the gating mechanism rather than to architectural or training discrepancies.
@@ Results.
Table 3 reports mean accuracy over five seeds. DGT improves over both baselines across all datasets. On the simpler sentiment tasks (Rotten Tomatoes, IMDB) and the four-class AG News, gains are modest (typically +1-2%), reflecting that these benchmarks are near-saturated with strong baselines. Importantly, DGT maintains or improves performance across all expansion settings, indicating that input-dependent gating introduces no regressions even in low-complexity regimes. The advantages are most evident on 20 Newsgroups, where DGT consistently surpasses both baselines, demonstrating its capacity to handle fine-grained, high-variance classification with greater robustness. On MNLI, DGT consistently outperforms both baselines across expansion factors, with improvements of up to +3% over the vanilla Transformer and +1-1.5% over DT, underscoring its effectiveness in large-scale natural language inference. Overall, the findings indicate that input-dependent gating yields consistent improvements on easier benchmarks and delivers pronounced advantages as task difficulty rises, underscoring its ability to scale reliably and adapt to diverse linguistic settings.
@@
Figure 2: Attention-rollout heatmaps for DGViT vs DViT vs ViT on CIFAR-100 (top two rows) and SVHN (bottom two rows). Bright yellow indicates high attention weights, whereas deep blue denotes low or suppressed attention. ViT disperses attention across both objects and background, DViT partially cancels common-mode noise but retains some distraction, while DGViT exhibits sharper boundary delineation and stronger alignment with task-relevant features.
@@
(b) Differential Transformer model attention
@@
Figure 3: Attention-rollout for a 20 Newsgroups sample. Bright yellow indicates strong positive attention, and deep blue denotes inhibitory weights. (a) The standard Transformer baseline spreads its attention diffusely across both function words and content tokens, failing to isolate the class‐defining terms. (b) The Differential Transformer partially suppresses common‐mode noise, but still assigns non‐negligible weight to other classes too. (c) Our DGT model sharply concentrates on the key parts of the input while attenuating filler words. The focused, input-dependent gating aligns attention with the correct comp.windows.x label and filters out background noise.
Vision (DGViT vs Baselines).
Fig. 2 illustrates attention-rollout heatmaps for DGViT, DViT, and ViT on CIFAR-100 and SVHN. While ViT spreads attention broadly, often assigning high weights to background regions alongside the object of interest, DViT applies a differential contrast mechanism that partly sharpens focus but still leaves residual noise. DGViT, by contrast, produces the most discriminative patterns: excitatory weights are concentrated on semantically salient structures (e.g., mushroom contours or digit strokes), while inhibitory weights actively suppress background clutter, resulting in clearer boundaries and a more reliable focus on task-relevant regions.
@@
Fig. 3 presents attention-rollout for (a) Transformer, (b) Differential Transformer, and (c) our DGT on a 20 Newsgroups sample, input text: "Most graphics systems I have seen have drawing routines that also specify a color for drawing, like Drawpoint(x,y,color) or Drawline(x1,y1,x2,y2,color) or Fillrectangle(x1,y1,x2,y2,color)…With X, I…", label: comp.windows.x. The vanilla Transformer disperses its weights broadly over both common function words and various candidate labels, failing to single out the true class while the Differential Transformer partly suppresses common‐mode noise yet still spreads attention to competing classes, diluting its focus on the correct one. DGT’s input-dependent gating, in contrast, concentrates focus on key class‐indicative words, filtering out irrelevant text and yielding a more interpretable, discriminative attention pattern, increasing attention mass on ground-truth and aligning sharply with comp.windows.x.
@@ 6 Discussion
Our results demonstrate that embedding an input-conditioned lateral inhibition mechanism into self-attention yields substantial gains in noise resilience and generalization across both vision and language tasks. M-DGSA consistently sharpens attention maps, assigning strong excitatory weights to semantically relevant regions while suppressing background noise, without requiring any task-specific tuning. Its capacity to delineate fine boundaries suggests applications in safety‐critical domains - such as lane‐marking detection in autonomous driving or edge delineation in medical imaging - where robust, fine‐grained attention is paramount. Despite these benefits, several limitations invite further exploration. We have so far evaluated primarily on mid- and large-scale benchmarks with synthetic noise, broadening the analysis to real-world distortions (e.g., motion blur, occlusions) may uncover additional strengths or weaknesses. Moreover, our study focused on self-attention, incorporating lateral inhibition into cross-attention - such as in vision-language models for VQA or image captioning - could enable dynamic, context-conditioned suppression of irrelevant inputs and further strengthen multimodal alignment. Looking forward, adapting lateral-inhibition gating to multimodal architectures, scaling up to natural corruption regimes, and developing hardware-efficient implementations that leverage the induced sparsity represent particularly promising directions.
@@ 7 Conclusion
We have introduced Multihead Differential Gated Self‐Attention (M‐DGSA), a lightweight yet powerful extension of multihead self‐attention that embeds per‐head input-dependent lateral inhibition into the dual‐softmax framework. Leveraging dynamic, context‐sensitive gating, M‐DGSA produces sharper and more selective attention maps, enhancing noise resilience and improving cross-domain performance on both vision and language benchmarks. Importantly, it integrates seamlessly into existing Transformer and ViT architectures with negligible overhead, making it directly applicable to a wide range of tasks.
Beyond empirical improvements, M‐DGSA advances the interpretability of attention mechanisms: its gating patterns mirror biological contrast‐enhancement, yielding clear visualizations of how models emphasize salient features and quench noise. We view this biologically inspired paradigm as a significant step toward more robust, trustworthy attention models for next‐generation deep learning.
Reproducibility Statement
We have taken several measures to facilitate reproducibility of our results. All datasets are standard public benchmarks with canonical train/test splits, requiring no additional preprocessing beyond what is provided in their official releases A.1, A.2. Hyperparameter settings, model sizes, and training schedules for both vision and language tasks are detailed in Tables 4-8 of the Appendix. Compute requirements and runtime estimates are reported in Appendix A.4, and additional ablation studies on initialization and gate depth are provided in Appendix A.5.
@@ A.1 Detailed description of the used datasets
MNLI. (Williams et al., 2018) The Multi-Genre Natural Language Inference corpus is a dataset of 433,000 sentence pairs annotated with textual entailment, genre and label.
AG News. (Del Corso et al., 2005) Derived from AG’s corpus of news articles, the AG News dataset comprises four categories - World, Sports, Business, and Science/Technology - by combining each article’s title and description and selecting the four largest classes. It contains $120,000$ training samples ($30,000$ per category) and $7,600$ test samples ($1,900$ per category).
20 Newsgroups. (Mitchell, 1997) 18,000~18,000 posts organized into 20 distinct newsgroups, each corresponding to a specific topic such as computer hardware, recreational activities, politics, science, or religion. We adopt the canonical split of $11,314$ training and $7,532$ test posts, yielding on average about $566$ training and $377$ test samples per class.
@@
Street View House Numbers (SVHN). (Netzer et al., 2011) A large-scale, real-world image dataset comprising over 600,000 32×32-pixel digit crops extracted from Google Street View house numbers. It includes 10 classes (digits 0-9): 73,257 training images, 26,032 test images, and 531,131 additional “extra” samples of comparatively easier digits. By capturing digits in varied natural scenes - with diverse backgrounds, lighting conditions, and occlusions - SVHN poses a substantially more challenging recognition task than MNIST.
ImageNet. (Deng et al., 2009) A large-scale, human-annotated image dataset organized according to the WordNet hierarchy, aiming to provide around 1000 images per concept, with tens of millions of labeled images overall. It was created to advance computer vision research by offering a high-quality benchmark for object categorization and supplying the large amounts of data needed for developing more generalizable machine learning methods.
@@ A.2 Dataset Splits
For each language dataset, AG News, IMDB, 20 Newsgroups, and Rotten Tomatoes, we further split the training fold into $80\%$ train and $20\%$ validation (stratified by class). All final results are reported on the official test set. For all vision benchmarks, CIFAR-10, CIFAR-100, Fashion-MNIST, and SVHN, we use the standard train/test partitions provided by each dataset and no further hold-out split was applied.
@@ A.3.1 Vision Benchmarks
For all vision datasets, we use the hyperparameters reported in Table 4. For ImageNet-1k, we set the batch size to 64 and the patch size to 16.
@@ A.3.2 Language Benchmarks
For AG News, IMDB, and 20 Newsgroups, all text‐classification experiments use the hyperparameters listed in Table 5. For Rotten Tomatoes, we instead employ 4 layers, a learning rate of $5\times 10^{-4}$, and 500 warmup steps. The hyperparameters for MNLI are reported separately in Table 6.
Table 5: Hyperparameters for small text datasets.
@@
Table 6: Hyperparameters for MNLI dataset.
Hyperparameter Value
Epochs 5
Batch size 128
Learning rate $3\times 10^{-4}$
Warmup steps 4000
Max sequence length 256
Minimum token frequency 2
Maximum vocabulary size 30,000
Embedding dimension 512
# layers 6
# heads 8
Dropout probability 0.1
Optimizer AdamW ($\beta_{1}=0.9,\;\beta_{2}=0.98,\;\epsilon\!=\!10^{-8}$)
Weight decay 0.01 (0 for bias & LayerNorm/RMSNorm weights)
LR scheduler Linear warmup (4000 steps) + linear decay
Gradient clipping Max-norm = 1.0
@@ A.4 Compute & Runtime
ImageNet-1k experiments ran on NVIDIA A100 (40 GB and 80 GB) GPUs. All other experiments ran in PyTorch on a single NVIDIA A10G (24 GB).
  • Vision: ${\sim}0.5$min/epoch on CIFAR-10/100, ${\sim}0.6$ min on FashionMNIST, ${\sim}0.76$ min on SVHN, and ${\sim}31$min/epoch on ImageNet.
  • Language: ${\sim}0.3$min/epoch on Rotten Tomatoes, ${\sim}0.39$ min on IMDB, ${\sim}0.57$ min on AG News, ${\sim}0.5$ min on 20NewsGroups, and ${\sim}18$min/epoch on MNLI.
With DT and DGT, our goal was to avoid increasing - and in some cases even to reduce - the number of learnable parameters relative to the standard Transformer. This reduction stems from the use of SwiGLU, which splits projections in half, whereas the Transformer employs GeLU over the full hidden dimension.
The standard ViT has 5.9M parameters across CIFAR-10, CIFAR-100, FashionMNIST, and SVHN. For DViT and DGViT, the parameter count varies with the projection factor: a $\times 2$ projection yields 3.8M parameters, while a $\times 4$ projection increases this to 5.4M.
In the ImageNet experiments, all models maintain a comparable computational cost of 1.2 GFLOPs regardless of architecture. The baseline ViT has 5.7M learnable parameters and uses 2.7 GB of memory. DViT and DGViT with a $\times 4$ projection keep the parameter count at 5.7M, but memory usage rises to 4.3 GB. The $\times 16/3$ projection further increases parameters to 6.7M, while leaving the memory footprint unchanged. These results indicate that our architectural modifications primarily affect parameter count and memory requirements, while leaving overall FLOPs constant.
The number of trainable parameters in the language tasks of Rotten Tomatoes, IMDB, AGNews and 20 Newsgroups are reported in Table 7 and for the MNLI in Table 8.
Table 7: Number of learnable parameters for language benchmarks.
@@
Transformer 5.15.1M 27.627.6M 15.115.1M 13.313.3M
DT/DGT
   proj. $\times 2$ 3.83.8M 25.725.7M 13.213.2M 1212M
   proj. $\times 4$ 4.64.6M 26.926.9M 14.414.4M 12.812.8M
   proj. $\times 16/3$ 5.15.1M 27.727.7M 15.215.2M 13.413.4M
Table 8: Metrics from the MNLI experiments.
No. of learnable params Memory usage (in MB) GFLOPs
Transformer 31.4M 1.3k 4.4
DT/DGT
   proj. $\times 2$ 26.7M 2.1k 3.2
   proj. $\times 4$ 31.4M 2.1k 4.4
   proj. $\times 16/3$ 34.6M 2.1k 5.2